4. 04 points I Previous Answers If 27.4 mL of silver nitrate solution reacts with excess potassium chioride solution to yield 0889 g of precipitate, what is the molarity of silver ion in the original solution?
media%2F86e%2F86e5fb7a-3b08-446d-8380-65
Answer #1

Answer 1.

The chemical reaction occuring is

AgNO_3 + KCl \leftrightarrow AgCl + KNO_3

Molar mass of AgCl = 143.5 g/mol

Therefore,

moles = mass/molar mass = 0.899/143.5 = 0.00627 moles

Therefore,

Moles of AgCl = Moles of AgNO3 = Moles of Ag+

Hence,

Molarity of silver ion = moles/volume = 0.00627/0.0274 = 0.23 M

Answer 2.

The chemical reaction occuring is

FeBr_3 + K_3PO_4 \leftrightarrow FePO_4 + 3KBr

Formula = FePO4

Name = Iron(III) Phosphate